1/2:2 as 1/4:X

X=

Answer by jim_thompson5910(35256)

(1/2)/2 = (1/4)/x


(1/2)x = 2*(1/4)


x/2 = 2/4


4x = 2*2


4x = 4


x = 4/4


x = 1


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim